OpenStudy (anonymous):

Would any restrictions apply to the domain and range of those equations? Explain your reasoning using complete sentences.

OpenStudy (anonymous):

what are the equations?

OpenStudy (anonymous):

oh my sorry 2x-3y=-4 y-4=1(x-4) y=(-5/4)x+15.5

OpenStudy (anonymous):

@helpme1.2